OpenStudy (anonymous):

Find P( F | E ) E= Pointer lands on an even number F= Pointer lands on a number less than 4

OpenStudy (anonymous):



OpenStudy (anonymous):

can you tell from the table what \(P(E)\) is ?

OpenStudy (anonymous):

mmm not exactly, im not sure what do to with it

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

add

OpenStudy (anonymous):

they equal 1

OpenStudy (anonymous):

there are two even numbers \(2\) and \(4\)

OpenStudy (anonymous):

i didn't mean add all the numbers, they have to add to \(1\) i meant add the probabilities for getting an even number \[P(2)+P(4)\]

OpenStudy (anonymous):

ahh so .4?

OpenStudy (anonymous):

that is not the final answer, but you need that number, yes

OpenStudy (anonymous):

\[P(F|E)=\frac{P(F\cap E)}{P(E)}\] so we need \(P(E)=.4\) as out denominator

OpenStudy (anonymous):

so now find P(F)?

OpenStudy (anonymous):

as * our denominator now that we know it is \(.4\) you need \(P(E\cap F)\)

OpenStudy (anonymous):

you need \(P(E\cap F)\) not \(P(F)\)

OpenStudy (anonymous):

oh okay, so how do i find that?

OpenStudy (anonymous):

in order to find it, you have to understand what \(E\cap F\) means in simple english then i will be obvious

OpenStudy (anonymous):

E= Pointer lands on an even number F= Pointer lands on a number less than 4 join these two statements with the word "AND" and then restate it in plain english

OpenStudy (anonymous):

if it is not clear let me know

OpenStudy (anonymous):

oh so .2?

OpenStudy (anonymous):

cause 2 would be the only even number right?

OpenStudy (anonymous):

well \(E\cap F\) is a set, it means "you roll an even number And a number less that 4" in english it means "your roll a two" but you are right \[P(E\cap F)=.2\]

OpenStudy (anonymous):

that completes finding the numbers you need for your answer

OpenStudy (anonymous):

you get the final answer?

OpenStudy (anonymous):

yes .5

OpenStudy (anonymous):

yeah \[\frac{.2}{.4}=\frac{2}{4}=\frac{1}{5}=.5\]